# Chrono[Uncaptioned image]: A Simple Blueprint for Representing Time in MLLMs

### v5

@@ Abstract
The recent success of Large Language Models (LLMs) has prompted the extension to the multimodal domain developing image-text Multimodal LLMs (MLLMs) and then video-text models. In this work, we investigate the challenge of contextual and temporal comprehension in video-language models by exploring the task of temporal localization in videos. To address this problem, prior works have developed complex task-specific architectures, novel modules to embed time into MLLMs, or leveraged additional input signals such as video transcripts to best encode contextual and temporal information. Interestingly, we find that most of these efforts are surpassed by a much simpler design. We introduce Chrono[Uncaptioned image], a universal sequence blueprint that can be applied to an image-text pretrained MLLM. Through extensive ablations across different MLLM architectures, finetuning and zero-shot settings, and different datasets, we achieve a new SOTA in moment retrieval on the most widely used benchmarks Charades-STA, QVHighlights, ActivityNet Captions, and grounded video question answering on NeXT-GQA. [^4] Our code is available under https://github.com/sudo-Boris/mr-Blip.
Machine Learning, ICML
@@ 1 Introduction
The recent success of pretrained large language models (LLMs) (Brown et al., 2020; Zhang et al., 2022; Chung et al., 2022) has inspired the development of generative image-text pretrained multimodal large language models (MLLMs) (Alayrac et al., 2022; Li et al., 2023; Huang et al., 2024b) that can comprehend vision and language modalities jointly. However, due to higher computational and annotation costs, large-scale pretraining on video data is more challenging to scale. To circumvent the issue, recent studies leverage image-text pretrained models for image-to-video transfer learning (Yu et al., 2023, 2024; Lei et al., 2021b; Luo et al., 2022; Fang et al., 2021; Ju et al., 2022). Such models offer promising results in the direction of video-text retrieval (Luo et al., 2022), video captioning (Xu et al., 2023), or multiple choice video question answering (Yu et al., 2024). Yet, those tasks don’t require precise temporal understanding, whereas the task of moment retrieval (MR) requires the precise temporal localization of all moments associated with an open-ended natural-language query in an untrimmed video and has not been extensively explored.
@@
This task has multiple potential valuable applications, such as video search or video indexing, but most importantly, it can be leveraged as a testing bed for a concrete temporal understanding of video-language models. More concretely, it requires understanding, discrimination, and temporal localization of multiple events in potentially minutes-long videos given an open-ended natural-language query. In the context of MLLMs, it is an open question of how to best model this task as a sequence-to-sequence prediction task, specifically, how to encode time best so the model can reason about it and predict start and end timestamps correctly.
@@
Traditionally, prior works generally approach the challenge of moment retrieval by leveraging video features (Feichtenhofer et al., 2019) or CLIP (Radford et al., 2021) to train a complex, task-specific feature fusion module, which ultimately predicts a fixed set of candidates. As shown in Figure 1, recent and concurrent approaches employ MLLMs for temporally grounded video-language tasks, one of the most relevant being moment retrieval. However, they often require large-scale instruction tuning datasets (Ren et al., 2023; Qian et al., 2024; Yang et al., 2023), complex multi-stage training (Huang et al., 2024a), specialized architectures (Qian et al., 2024; Ren et al., 2023), or further input signal, such as video transcripts (Yang et al., 2023).
@@
In contrast, we develop Chrono[Uncaptioned image], a novel multi-modal input sequence blueprint that can enable image-text pretrained MLLMs to perform video moment retrieval. Interestingly, our thorough ablations demonstrate that our simple approach outperforms all prior methods, which add significantly more complexity, as shown in Figure 1. We show that this input design enables the best temporal grounding abilities both when finetuning image-text pretrained MLLMs or when using powerful MLLMs for zero-shot applications. Our approach achieves state-of-the-art (SOTA) results across the widely used benchmarks Charades-STA (Gao et al., 2017), QVHighlights (Lei et al., 2021a), and ActivityNet Captions (Krishna et al., 2017). We discuss the relevance of deliberate design choices by conducting extensive ablation studies on two MLLM backbones with different architectures, BLIP-2 (Li et al., 2023) and GPT-4o (OpenAI, 2024a). Besides the video moment retrieval datasets, we also consider NExT-GQA (Xiao et al., 2023), a dataset for grounded video QA (GVQA). The task there is to localize temporal segment(s) relevant to the question’s answer. Here, we assess Chrono’s ability to localize such video evidence without being trained for this task, as well as to generate an answer; again, we achieve SOTA results.
@@
Our main contributions are as follows: (i) We leverage image-text pretrained MLLMs to approach moment retrieval by casting it as an open-ended sequence-to-sequence problem. (ii) To enhance the temporal understanding of events in input videos, we design a novel multimodal input sequence and introduce Chrono[Uncaptioned image], a simple and universal blueprint for representing time in MLLMs. (iii) Chrono models improve the state-of-the-art on the widely used moment retrieval benchmarks (Gao et al., 2017; Lei et al., 2021a; Krishna et al., 2017) and achieves a new SOTA on grounded VQA (Xiao et al., 2023). (iv) Extensive experiments and ablations demonstrate the effectiveness of Chrono and its design choices, highlighting that deliberate exploration of simple methods can outperform complex, potentially over-engineered methods. (v) We make our code and models publicly available.
@@ 2 Related Work
2.1 Moment Retrieval Models
Moment Retrieval analyzes an untrimmed video and aims to find the relevant clip for a given open-ended natural language query. Conventional approaches fall into either proposal-based or proposal-free methods. Proposal-based methods learn to identify moment candidates given predefined proposals, e.g., sliding windows (Anne Hendricks et al., 2017; Gao et al., 2017) and temporal anchors (Chen et al., 2018; Wang et al., 2020), in a first stage. In a second stage, these candidates are further refined to better match the text query. Regression-based methods (Lei et al., 2021a; Yan et al., 2023; Liu et al., 2022; Mun et al., 2020; Zeng et al., 2020) are common among proposal-free approaches. These directly predict the temporal boundaries of a relevant moment, without needing a first stage to generate initial proposals. Nevertheless, proposal-free methods still predict a fixed number of candidate windows alongside a confidence score, based on which the predictions are sorted. Using a fixed number of proposals has some downsides, like not generalizing to videos with a higher number of actions than the data used for tuning the number of proposals hyperparameter or wasted computation on duplicate proposals.
Recent works have explored various approaches to temporal modeling using multimodal LLMs. Ren et al. (2023) introduce time-sensitive frame embeddings by incorporating temporal information through Q-Former, into their model, TimeChat. Huang et al. (2024a) train VTimeLLM in a three-stage training setup, including extensive video-language pretraining. Momentor (Qian et al., 2024) proposes a dedicated Temporal Perception Module, and Lita (Huang et al., 2025) explores special temporal tokens. While these recent and in various cases concurrent approaches demonstrate the potential of MLLMs for temporal understanding, they often require large-scale instruction tuning datasets (TimeChat, Momentor), complex multi-stage training (VTimeLLM), or specialized architectures (Momentor). In contrast, our work shows that careful design choices in representing temporal information can enable moment retrieval without any prior training on the moment retrieval task. Furthermore, when combined with direct finetuning on downstream tasks, our Chrono[Uncaptioned image] blueprint can achieve superior performance without the need for expensive pretraining or architectural complexity. Through systematic ablations, we demonstrate that simple design choices - such as using absolute integer timestamps - can outperform more complex approaches that introduce special tokens or dedicated temporal modules.
2.2 Image-to-Video Transfer Learning
@@
The limited public availability of high-quality large-scale video-language datasets, together with the immense compute requirements of training using videos, pose a significant challenge for large-scale video-language pretraining. The idea of leveraging image-language pretrained models for image-to-video transfer learning by utilizing a limited number of video frames to enhance learning efficiency has proven effective as evidenced by many recent publications (Yu et al., 2023, 2024; Cao et al., 2022; Fang et al., 2021, 2022; Ju et al., 2022; Lei et al., 2021b; Luo et al., 2022; Ma et al., 2022; Lei et al., 2022; Xue et al., 2023; Wang et al., 2022). In particular, (Yu et al., 2024) makes use of the contextual understanding of MLLMs to perform video QA tasks. The authors train a BLIP-2 model to individually select 4 question-relevant frames and, using those, finetune another BLIP-2 to answer the question. We also leverage the pretrained BLIP-2 model and finetune it on the downstream task of moment retrieval, achieving big localization improvements over Yu et al. (2024). Furthermore, we show that strong image-text models like GPT-4o (OpenAI, 2024b), which cannot perform moment retrieval natively, achieve competitive results using our Chrono blueprint.
3 Chrono[Uncaptioned image]
@@
Figure 2: ChronoRefer to caption model overview. In this work, we demonstrate that our framework is model-agnostic and can adapt an image-text pretrained MLLM even in a zero-shot (no finetuning) setting. We construct the prompt for the MLLM by interleaving the frame embeddings and the timestamps of each sampled frame, followed by the video duration, the moment retrieval query, and a task prompt (the latter being not visualized.) The MLLM outputs a sequence of potentially multiple retrieved moments by predicting the global BOS and EOS tokens, denoted by ##\##, the start of window and end of window tokens denoted as #$\#_{start}$ and #$\#_{end}$, and the respective start and end times for each window. In the case of finetuning a model (Chrono-BLIP), we freeze the MLLM and only finetune additional adapter layers leveraging parameter-efficient finetuning (Hu et al., 2022).
@@
To explore how to best represent time in an MLLM, we select the task of moment retrieval as a suitable test bed. The task is to temporally localize all relevant moments in an untrimmed video given an open-ended natural language query. Therefore, a key challenge is to effectively model the contextual and temporal relationship between the different events in the video, for example, as illustrated in Figure 2, a man is talking to the camera followed by him watching a show. Furthermore, due to the amount of information in a video, it is computationally unfeasible to leverage all frames of a video as context in a single forward pass. To tackle these challenges, we cast moment retrieval as a language modeling task and leverage the contextual understanding ability of generative MLLMs to interpret and comprehend the semantic and temporal action development in frames that represent a video. We first thoroughly explore the design space of timestamps in MLLMs. Our analysis resulted in our proposed recipe, Chrono, that can enable image-text pretrained MLLMs to interpret video frames and their respective temporal grounding in both a finetuning and even zero-shot setting. In Section 3.1, we present our proposed blueprint, Chrono. In Section 3.2 we discuss how our model-agnostic Chrono blueprint can be leveraged for adapting an MLLM using finetuning and a zero-shot setup.
@@ 3.1 Chrono blueprint
Our work investigates how to enable image-text pretrained MLLMs to comprehend video and time for temporal grounding tasks in videos given a natural language query. To do so, we cast the traditional moment retrieval task as an open-ended sequence-to-sequence problem, where we design a novel multimodal input sequence that contains (a) the visual semantic context given the sampled frames of the video, (b) the temporal context for each event, which we model through by the temporal location of each frame (i.e., a timestamp) as well as the total video duration, and (c) the query in the form of a natural language description. The output of our model is a time interval sequence where each time interval represents a moment associated with the query and follows the formatting of a nested list of (potentially) multiple moments with a start and end time. Notably, we do not add any new special tokens and use the vocabulary native to the backbone MLLM, which allows our framework to be more general and applicable in a zero-shot setting.
To model contextual and temporal relationships between the different actions in a video and a natural language query, we cast moment retrieval as a sequence-to-sequence task and develop a novel multimodal input sequence. As illustrated in Figure 2, it consists of frames $f_{n}$, timestamps $t_{n}$ ($n=1,...,F$), video duration $d$, query $q$, and task prompt $p$ stating the task at hand and, in the zero-shot setting, a format-adherence prompt, which we shown in Appendix Figure 4. Notably, there are multiple ways of representing timestamps $t_{n}$. Time can be represented in absolute or relative form, e.g., “79.9” seconds or “0.40” for a relative position (e.g. 0 being the start of the video and 1 the end); it could be a decimal number or be rounded to the nearest full integer value, e.g., “80” seconds or “40” for a relative position; the timestamps could be concatenated after the frames $f_{n}$ or interleaved with them. As illustrated in Figure 2, we find that representing time as simple text tokens of whole seconds (integer absolute form) and interleaving them with corresponding frames yields the best performance. This finding leads to our final design for our multimodal input sequence

### v6

# Chrono: A Simple Blueprint for Representing Time in Multimodal Large Language Models

@@ Abstract
The recent success of Large Language Models (LLMs) has prompted the extension to the multimodal domain, developing image-text Multimodal LLMs (MLLMs) and then video-text models. In this work, we investigate the challenge of contextual and temporal comprehension in video-language models by exploring the task of temporal localization in videos. To address this problem, prior works have developed complex task-specific architectures, novel modules to embed time into MLLMs, or leveraged additional input signals such as video transcripts to best encode contextual and temporal information. We find that most of these efforts are surpassed by a much simpler design. We introduce Chrono[Uncaptioned image], a universal sequence blueprint that can be applied to any image-text pretrained MLLM. In extensive experiments spanning different MLLM architectures and sizes, finetuning and zero-shot settings, we demonstrate new state-of-the-art results in moment retrieval on the widely used benchmarks Charades-STA, QVHighlights, and ActivityNet Captions, as well as in grounded video question answering on NExT-GQA.  [^1] The code is available at https://github.com/sudo-Boris/mr-Blip.
@@ 1 Introduction
The recent success of pretrained large language models (LLMs) [brown2020gpt3, zhang2022opt, chung2022t5] has inspired the development of generative image-text pretrained multimodal large language models (MLLMs) [alayrac2022flamingo, li2023blip2, huang2023kosmos1] that can comprehend vision and language modalities jointly. However, due to higher computational and annotation costs, large-scale pretraining on video data is more demanding. To circumvent the issue, recent studies leverage image-text pretrained models for image-to-video transfer learning [yu2023EILEV, yu2023SeViLA, lei2021CLIPBERT, luo2021clip4clip, fang2021clip2video, ju2022prompting]. Such models offer promising results in the direction of video-text retrieval [luo2021clip4clip], video captioning [xu2023mplug], or multiple choice video question answering [yu2023SeViLA]. Yet, those tasks don’t require precise temporal understanding, whereas the task of moment retrieval (MR) requires the precise temporal localization of all moments associated with an open-ended natural-language query in an untrimmed video. This ability has not been extensively explored for such models.
@@
The moment retrieval task has multiple valuable applications, such as video search or video indexing. Additionally, we highlight temporal grounding as a test bed for temporal understanding of video-language models. More concretely, it requires understanding, discrimination, and temporal localization of multiple events in potentially minutes-long videos given an open-ended natural-language query. In the context of MLLMs, it is an open question of how to best model this task as a sequence-to-sequence prediction task. Specifically, how to encode time best so the model can reason about it and predict start and end timestamps correctly.
@@
Traditionally, prior works generally approach the challenge of moment retrieval by leveraging video features [feichtenhofer2019slowfast] or CLIP [radford2021CLIP] to train a complex, task-specific feature fusion module, which ultimately predicts a fixed set of candidate windows. As shown in Figure 1, recent approaches employ MLLMs for temporally grounded video-language tasks, one of the most relevant being moment retrieval. However, they often require large-scale instruction tuning datasets [Ren2023TimeChatAT, qian2024momentor, yang2023vid2seq], complex multi-stage training [huang2024vtimellm], specialized architectures [qian2024momentor, Ren2023TimeChatAT], or further input signal, such as video transcripts [yang2023vid2seq].
@@
In contrast, we develop Chrono[Uncaptioned image], a novel multimodal input sequence blueprint that enables image-text pretrained MLLMs to better understand time and achieve higher quality video moment retrieval. Chrono[Uncaptioned image] interleaves language timestamps with video frames, and appends duration information. By using the native language input space of the model, it avoids adding special tokens or any architectural modification. Extensive ablations demonstrate that this simple approach outperforms all prior methods, which add significantly more complexity, as shown in Figure 1. We show that this input design enables the best temporal grounding abilities both when finetuning image-text pretrained MLLMs or when using larger MLLMs in a zero-shot manner. Chrono[Uncaptioned image] achieves state-of-the-art (SOTA) results across the widely used benchmarks Charades-STA [gao2017charadesSTA], QVHighlights [lei2021qvhighlights], and ActivityNet Captions [krishna2017ActivityNetCaptions]. We discuss the relevance of deliberate design choices by conducting extensive ablation studies on MLLMs with different architectures, BLIP-2 [li2023blip2], Qwen2.5-VL [Qwen2.5-VL] and GPT-4o [GPT4o]. Besides the video moment retrieval datasets, we also consider NExT-GQA [Xiao2023NExT-QGA], a dataset for grounded video QA (GVQA). The task there is to localize temporal segment(s) relevant to the question’s answer. Here, we assess Chrono’s ability to localize such video evidence without being trained for this task, as well as to generate an answer. Again, we show constitent improvements over prior approaches which are significantly more complex.
@@
The main contributions are as follows: (i) We leverage image-text pretrained MLLMs to approach moment retrieval by casting it as an open-ended sequence-to-sequence problem. (ii) To enhance the temporal understanding of events in input videos, we design a novel multimodal input sequence and introduce Chrono[Uncaptioned image], a simple and universal blueprint for representing time in MLLMs. (iii) Chrono models, derived from BLIP-2, GPT-4o and Qwen2.5-VL, improve the state-of-the-art on the widely used moment retrieval benchmarks [gao2017charadesSTA, lei2021qvhighlights, krishna2017ActivityNetCaptions] and achieves a new state-of-the-art on grounded VQA [Xiao2023NExT-QGA]. (iv) Extensive experiments and ablations demonstrate the effectiveness of Chrono and its design choices, highlighting that deliberate exploration of simple methods can outperform complex, potentially over-engineered methods. (v) We make the code and models publicly available.
Since the release of our early arXiv preprint [meinardus2024surprising], Chrono has been adopted by several works to improve temporal understanding in moment retrieval, question answering, and captioning. More details are provided in Section 2. An early version of this work appeared at the ICCV 2025 workshop on Multimodal Representation and Retrieval [Meinardus_2025_ICCV].
@@ 2 Related Work
Image-to-Video Transfer Learning. The limited public availability of high-quality large-scale video-language datasets, together with the immense compute requirements of training using videos, pose a significant challenge for large-scale video-language pretraining. The idea of leveraging image-language pretrained models for image-to-video transfer learning by utilizing a limited number of video frames to enhance learning efficiency has proven effective as evidenced by many recent publications [yu2023EILEV, yu2023SeViLA, cao2022locvtp, fang2021clip2video, Fang2022TransferimImageCLIP, ju2022prompting, lei2021CLIPBERT, luo2021clip4clip, ma2022xclip, lei2022revealing, xue2023clipvip, wang2022objectaware].
In particular, [yu2023SeViLA] makes use of the contextual understanding of MLLMs to perform video QA tasks. The authors train a BLIP-2 model to individually select 4 question-relevant frames and, using those, fine-tune another BLIP-2 to answer the question.
We also leverage the pretrained BLIP-2 model and finetune it on the downstream task of moment retrieval, achieving significant localization improvements over yu2023SeViLA. Furthermore, we show that strong image-text models like GPT-4o [openai2024gpt4technicalreport], which cannot perform moment retrieval natively, achieve competitive results using the Chrono blueprint.
Moment Retrieval Models. Moment Retrieval (MR) tasks requires to analyze a video and find the relevant clip given an open-ended natural language query. Conventional approaches fall into either proposal-based or proposal-free methods. Proposal-based methods learn to identify moment candidates given predefined proposals, e.g., sliding windows [hendricks2017localizing, gao2017charadesSTA] and temporal anchors [chen2018temporalgrounding, wang2019temporally], in a first stage. In a second stage, these candidates are further refined to better match the text query. Regression-based methods [lei2021qvhighlights, yan2023unloc, liu2022umt, mun2020localglobal, zeng2020dense] are common among proposal-free approaches. These directly predict the temporal boundaries of a relevant moment, without needing a first stage to generate initial proposals. Nevertheless, proposal-free methods still predict a fixed number of candidate windows alongside a confidence score, based on which the predictions are sorted. Using a fixed number of proposals has some downsides, like not generalizing to videos with a higher number of actions than the data used for tuning the number of proposals hyperparameter or wasted computation on duplicate proposals.
MLLM-based temporal grounding. With the advent of multimodal large language models (MLLMs), several works have approached MR by using multimodal LLMs, leveraging their visual capabilities and extensive world knowledge. However, powerful MLLMs still struggle to understand and express time precisely, which impedes their direct use for MR. Therefore, Ren2023TimeChatAT, huang2024vtimellm, qian2024momentor, huang2025lita focus on how to represent time inside MLLMs. Ren2023TimeChatAT introduce time-sensitive frame embeddings by incorporating temporal information through Q-Former, into their model, TimeChat. huang2024vtimellm train VTimeLLM three stages, including extensive multiple-event video-language pretraining and high-quality video-instruction tuning. qian2024momentor present Momentor, which uses a dedicated Temporal Perception Module. huang2025lita explore the use of special temporal tokens for LITA. These works, require large-scale instruction-tuning datasets (TimeChat [Ren2023TimeChatAT], Momentor [qian2024momentor]), complex multi-stage training (VTimeLLM [huang2024vtimellm],LITA huang2025lita), or specialized architectures (Momentor [qian2024momentor],LITA huang2025lita). Distinct from that, this work shows that careful design choices in representing temporal information can enable moment retrieval using off the shelf MLLMs without any architectural modifications. Furthermore, when combined with direct finetuning on downstream tasks, the Chrono[Uncaptioned image] blueprint introduced in this work can achieve superior performance without the need for expensive pretraining or architectural complexity. Through systematic ablations, we demonstrate that simple design choices - such as using absolute integer timestamps - can outperform more complex approaches that introduce special tokens or dedicated temporal modules. Alternative, recent work has shown that timestamping beyond language, i.e. frames directly in visual input space, is useful when using models with more powerful visual capabilities numberit2025.
Influence of the Chrono[Uncaptioned image] blueprint on subsequent work. Inspired by the early arXiv preprint version of this work [meinardus2024surprising], several recent methods employ the Chrono[Uncaptioned image] blueprint to achieve strong performance on moment retrieval tasks. lu2024llavamrlargelanguageandvisionassistant extend the LLaVA vision-language model for moment retrieval by using Chrono[Uncaptioned image] blueprint, in addition to frame selection and token compression to process longer videos. Similarly, unitime2025 employ the Chrono[Uncaptioned image] blueprint, together with an adaptive per frame token rate and hierarchical segment prediction at inference, to better handle the moment retrieval task in long videos. zeng2025tempo further develop the Chrono[Uncaptioned image] blueprint to employ a fixed number of tokens per timestamp, and use RL to teach the model to perform moment retrieval in a more flexible manner.
Besides moment retrieval, time representation is also relevant for multiple video tasks. The Chrono[Uncaptioned image] blueprint has also been used to improve temporal understanding in a variety of video tasks. dtos2025 employ Chrono[Uncaptioned image] to obtain better temporal representation, which proves crucial in the Referring video Object Segmentation task.
Additionally, videollama3 adopt a similar timestamping scheme, demonstrating improved performance across diverse video QA settings, including general and long-video understanding, temporal reasoning, and temporal grounding.
The impact of Chrono[Uncaptioned image] goes beyond influencing the way time is instilled in video inputs, as its SOTA performance as a moment retrieval model is used by zeng2025distimedistributionbasedtimerepresentation to pseudo-label the segments corresponding to video captions. The resulting model demonstrates enhanced dense video captioning and video question answering, in addition to moment retrieval.
3 Chrono[Uncaptioned image]: A Blueprint for Time Representation in MLLMs
@@
Figure 2: ChronoRefer to caption model overview. The ChronoRefer to caption blueprint is model-agnostic and can adapt an image-text pretrained MLLM even in a zero-shot (no finetuning) setting. We construct the input for the MLLM by interleaving the frame embeddings and the timestamps of each sampled frame, followed by the video duration, the moment retrieval query, and a task prompt (the latter being not visualized.) The MLLM outputs a sequence of potentially multiple retrieved moments by predicting the global BOS and EOS tokens, denoted by #\#, the start of window and end of window tokens denoted as $_{start}$ and $_{end}$, and the respective start and end times for each window. In the case of finetuning a model (e.g. Chrono-BLIP or Chrono-Qwen), we freeze the MLLM and only finetune additional adapter layers leveraging parameter-efficient finetuning [hu2021lora].
@@
To explore how to best represent time in an MLLM, we first select the task of moment retrieval as a suitable test bed. The task is to temporally localize all relevant moments in an untrimmed video given an open-ended natural language query. Therefore, a key challenge is to effectively model the contextual and temporal relationship between the different events in the video, for example, as illustrated in Figure 2, a man is talking to the camera followed by him watching a show. Furthermore, due to the amount of information in a video, it is computationally unfeasible to leverage all frames of a video as context in a single forward pass. To tackle these challenges, we cast moment retrieval as a language modeling task and leverage the contextual understanding ability of generative MLLMs to interpret and comprehend the semantic and temporal action development in frames that represent a video. We first thoroughly explore the design space of timestamps in MLLMs. This analysis resulted in the proposed recipe, Chrono, that can enable image-text pretrained MLLMs to interpret video frames and their respective temporal grounding in both a finetuning and even zero-shot setting. In Section 3.1, we present the proposed blueprint, Chrono. In Section 3.2 we discuss how the model-agnostic Chrono blueprint can be leveraged for adapting an MLLM using finetuning and a zero-shot setup.
@@ 3.1 Chrono blueprint
This work investigates how to enable image-text pretrained MLLMs to comprehend video and time for temporal grounding tasks in videos given a natural language query. To do so, we cast the traditional moment retrieval task as an open-ended sequence-to-sequence problem, where we design a novel multimodal input sequence that contains (a) the visual semantic context given the sampled frames of the video, (b) the temporal context for each event, which we model using the temporal location of each frame (i.e., the timestamp) as well as the total video duration, and (c) the query in the form of a natural language description. The model outputs a time interval sequence where each time interval represents a moment associated with the query and follows the formatting of a nested list of (potentially) multiple moments with a start and end time. Notably, we do not add any new special tokens and use the vocabulary native to the backbone MLLM, which makes the framework more general and applicable in a zero-shot setting.
To model contextual and temporal relationships between the different actions in a video and a natural language query, we cast moment retrieval as a sequence-to-sequence task and develop a novel multimodal input sequence. As illustrated in Figure 2, it consists of frames $f_{n}$, timestamps $t_{n}$ (n=1,,Fn=1,...,F), video duration $d$, query $q$, and task prompt $p$ stating the task at hand and, in the zero-shot setting, a format-adherence prompt, which we show in Appendix Figure 4. Notably, there are multiple ways of representing timestamps $t_{n}$. Time can be represented in absolute or relative form, e.g., “79.9” seconds or “0.40” for a relative position (e.g. 0 being the start of the video and 1 the end); it could be a decimal number or be rounded to the nearest full integer value, e.g., “80” seconds or “40” for a relative position; the timestamps could be concatenated after the frames $f_{n}$ or interleaved with them. As illustrated in Figure 2, we find that representing time as simple text tokens of whole seconds (integer absolute form) and interleaving them with corresponding frames yields the best performance. This finding leads to the final design for the multimodal input sequence